Every letter written R277 Every letter received R123 Drafting of summons R1044 Drafting of an affidavit R69, 50 per page or any part thereof Telephonic consultations R147,50 per 15 minutes or any part thereof Consultation R948,70 per 30 minutes or any part thereof Stamps R9.40 VAT excl VAT @ 15% During the past month Mahlobogwane Attorneys rendered the following services: Received 12 letters Wrote and posted 11 letters Drafted 3 summonses Drafted three affidavits1 consisting of 6 pages, 1 consisting of 3½ pages and 1 consisting of 6½ pages Consulted twice telephonically first for 28 minutes and later 19 minutes Consultation twice: first for 33 minutes and later for 1 hour and 32 minutes Instructed an Advocate, Mac Mafarani, to draft a document. The Advocate charged R3660,50 VAT excl Client went to Makankane MediClinic to consult a doctor, and source out a medico-legal report. The doctor charged R7 780.50 VAT excl. Draft a comprehensive statement of account for Mahlobogwane Attorneys to MacBuff Phosa


1
Expert's answer
2022-02-28T17:07:03-0500

letters received =12*123=1476

written=11*277

drafted summonses=3*1044 =3132

drafted affidavit=1*69.50*6

3.5*69.50 =243.25

6.5*69.50 =451.75

consultation telephonic=147.50 for 15min

1for 28min =275.33

2for19min =186.83

consultation =948.70 for 30 min

1 for 33min =1043.57

2 for 92 min =2909.34

advocate charge =3660.50 excl vat

mediclinic charge =7780.50 excl vat